OpenStudy (anonymous):

What is the slope of the line passing through the points (–1, 3) and (4, –7)? A. 2 B. 3/4 C. -4/3 D. -2

OpenStudy (smilie2015):

with the help of slope formula \[\frac{{y2 - y1}}{{x2 - x1}}\] you can find as x1 =-1 ,y1 = 3,x2=4, y2= -7 put these values in formula you will get the slope

OpenStudy (smilie2015):

m=-7-3/4-(-1) =-10/5 =-2

OpenStudy (smilie2015):

answer is D

OpenStudy (anonymous):

tysm